OpenStudy (anonymous):

find the slope if it exists, of the line containing the pair of points 5,8 and 9,-4

OpenStudy (anonymous):

slope = y2 - y1 / x2-x1 = 8 - (-4) / 5 - 9 = 12 / -4 = -3

OpenStudy (anonymous):

do you follow that lizz?

OpenStudy (anonymous):

sort of